# A Unified Framework for Iris Anti-Spoofing: Introducing IrisGeneral Dataset and Masked-MoE Method

### v1

@@ Introduction
Two main contributions are proposed in our MMoE method: Mask Mechanism, and Cosine Distance Loss. To prove the effectiveness of our contributions, we provide the ablation study on Protocol-1 and Protocol-2. Among that, the CLIP means the vanilla CLIP framework, while S is SoftMoE, L is the cosine distance between experts, and M is the Mask Mechanism. So for example, CLIP+S+M+L means CLIP framework with SoftMoE, Mask Mechanism, and cosine distance between experts, it is briefly called MMoE with CLIP (Ours). The evaluation results are shown in Table 6, and the ACERs’ trends are demonstrated in Figure 6. It could be observed that after adding the MoE ”S” into the vanilla CLIP, the performance significantly declined, especially on Protocol-2, which may caused by the overfitting problem of introducing MoE. After adding our contributions L and M, the performances improved on both Protocols. Specifically, to compare the ablation results, the CLIP+S+L is better on Protocol-2 which focuses on generalization, while the CLIP+S+M is better on Protocol-1 which focuses on average performance. That proves our contributions Cosine Distance Loss ”L” could improve the model’s generalization ability and Mask Mechanism ”M” significantly alleviates the overfitting problem caused by MoE. The CLIP+S+M+L (Ours) shows the best performance among these combined models, the ACER is 1.23% on Protocol-1 and 7.59% on Protocol-2. That good performance is thanks to both alleviating overfitting and improving generalization with our two contributions.
@@
Table 6: The evaluation results of each ablation on Protocol-1 and Protocol-2.
Figure 6: The ACER results of each ablation. It could be found that the addition of MoE ”S” makes the model worse, which may be caused by overfitting. Our contributions L and M significantly alleviate this problem and effectively improve the generalization performance.
Conclusion and Future Work
In this paper, we proposed a unified framework for iris anti-spoofing, which includes an IrisGeneral dataset and the Mased-MoE method. It significantly improves the generalization of iris anti-spoofing models and alleviates overfitting problems. In future work, we will focus on studying the boundary of the MMoE’s adaptive capability.

### v2

@@ V-D Ablation Study
In ablation study part, we evaluate the effectiveness of the components of our MMoE model. Two main contributions are proposed in our MMoE method: Mask Mechanism, and Cosine Distance Loss. To practically confirm the effectiveness of our contributions, we provide the ablation study on Protocol-1 and Protocol-2. In this part, the CLIP means the vanilla CLIP framework, S is SoftMoE, L is the cosine distance between experts, and M is the Mask Mechanism. For example, CLIP+S+M+L means CLIP framework with SoftMoE, Mask Mechanism, and cosine distance between experts, which is briefly called MMoE with CLIP or Ours. The evaluation results are shown in Table V, and the trends of ACERs, ACCs, AUCs and EERs are demonstrated in Figure 8. It could be observed that after adding the Soft MoE ”S” into the vanilla CLIP, the performance significantly declined, especially on Protocol-2, which might be caused by the overfitting problem of introducing MoE. After adding our contributions L and M, the performances improved on both Protocols. Specifically, to compare the ablation results, the CLIP+S+L is better on Protocol-2 which focuses on generalization, while the CLIP+S+M is better on Protocol-1 which focuses on average performance. That proves our contributions Cosine Distance Loss ”L” could improve the model’s generalization ability and Mask Mechanism ”M” significantly alleviates the overfitting problem caused by MoE. The CLIP+S+M+L (Ours) shows the best performance among these combined models, the ACER is 1.23% on Protocol-1 and 7.59% on Protocol-2. That good performance is thanks to both alleviating overfitting and improving generalization with our two contributions.
Consequently, the experimental results show that the Maks Mechanism could truly reduce the overfitting caused by the MoE, the Cosine Distance Loss could truly improve the generalization of the whole model, and both of these two components could improve the average performance of the detection of spoofing iris.
@@
TABLE V: The evaluation results of each ablation on Protocol-1 and Protocol-2.
VI Conclusion and Future Work
In this paper, we proposed a unified framework for iris anti-spoofing, which includes an Iris Anti-Spoofing Cross-Domain-Testing (IAS-CDT) Protocol and a Masked-MoE (MMoE) method. The IAS-CDT Protocol provides an evaluation standard for testing the generalization of iris recognition models. As both the theoretical analysis and the experimental confirmations demonstrates, our MMoE method could significantly improves the generalization of iris anti-spoofing models and alleviates overfitting problems. In future work, we will focus on studying the boundary of the MMoE’s adaptive capability.